OpenStudy (anonymous):

The measure of the supplement of an angle is 12 more than twice the measure of the angle. Find the measures of the angle and its supplement.

OpenStudy (anonymous):

@bobwhiz why did you delete your answer?

jimthompson5910 (jim_thompson5910):

let x = the angle so 180 - x is the supplement

jimthompson5910 (jim_thompson5910):

because "The measure of the supplement of an angle is 12 more than twice the measure of the angle", we know that 180 - x = 2x + 12

jimthompson5910 (jim_thompson5910):

solve that for x and tell me what you get

OpenStudy (anonymous):

I do not now how to solve for X

jimthompson5910 (jim_thompson5910):

180 - x = 2x + 12 180 - 12 = 2x + x 168 = 3x x = ??

OpenStudy (anonymous):

x = 56?

jimthompson5910 (jim_thompson5910):

so if x = 56, then what is 180 - x

OpenStudy (anonymous):

124?

jimthompson5910 (jim_thompson5910):

good, the original angle is 56 degrees the supplement is 124 degrees

OpenStudy (anonymous):

when you did "180 - x = 2x + 12 180 - 12 = 2x + x" how did you know how to do that?

OpenStudy (anonymous):

like how to you simplify

jimthompson5910 (jim_thompson5910):

180 - x is the supplement

jimthompson5910 (jim_thompson5910):

oh i see

jimthompson5910 (jim_thompson5910):

I added x to both sides and I also subtracted 12 from both sides

OpenStudy (anonymous):

oh ok thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

wait is the supplement 180-x or 124 degrees?

jimthompson5910 (jim_thompson5910):

180-x is the supplement of x 124 is the supplement of 56

jimthompson5910 (jim_thompson5910):

the final answer they want is 124 for the supplement but you use 180 - x to get there

OpenStudy (anonymous):

ohh thank you

jimthompson5910 (jim_thompson5910):

you're welcome